# Capturing Temporal Dynamics in Large-Scale Canopy Tree Height Estimation

### v1

@@ Acknowledgements
This work is supported via the AI4Forest project, which is funded by the German Federal Ministry of Education and Research (BMBF; grant number 01IS23025A) and the French National Research Agency (ANR). We gratefully acknowledge the substantial computational resources that were made available to us by the PALMA II cluster at the University of Münster (subsidized by the DFG; INST 211/667-1) as well as by the Zuse Institute Berlin. We also appreciate the hardware donation of an A100 Tensor Core GPU from Nvidia. We further gratefully acknowledge technical support from Google for this work.
@@
This study significantly advances forest monitoring by leveraging machine learning to incorporate temporal dynamics, resulting in the creation of tree canopy height maps over a four-year span (2019-2022) on a continental scale. Forests play a vital role in mitigating climate change by absorbing nearly half of human-generated carbon dioxide emissions (Friedlingstein et al., 2022). However, they are increasingly at risk due to climate change and human activities such as deforestation and land degradation (Anderegg et al., 2022). In support of global initiatives like the United Nations’ Sustainable Development Goals (SDGs)666https://sdgs.un.org/2030agenda, the Bonn Challenge777https://www.bonnchallenge.org, and the Glasgow Declaration888https://www.oneplanetnetwork.org/programmes/sustainable-tourism/glasgow-declaration, effective forest management and conservation are crucial for climate adaptation and mitigation. Our approach significantly benefits the estimation of above-ground biomass, a critical component among the 55 essential climate variables (ECVs)999https://gcos.wmo.int/site/global-climate-observing-system-gcos/essential-climate-variables/above-ground-biomass. A key improvement in this research is the enhanced detection of large trees, which is critical due to the power-law relationship between tree height and biomass. This advancement allows for more accurate biomass estimation, essential for assessing carbon stocks. Additionally, precise monitoring is crucial for validating carbon credit investments in the voluntary carbon market, ensuring the effectiveness and sustainability of forest growth projects while reducing potential leakage effects.
@@ Appendix A Data Handling Details
Handling big amounts of data is a challenge. To aid data handling and still having optimal projection accuracy, we downloaded, stored and preprocessed all data in the original Sentinel-2 tiling system. The tiling system is based on the UTM projections system but divides every UTM zone in tiles. Each UTM zone has its own projection101010For more details, we refer to: https://sentiwiki.copernicus.eu/web/s2-products.
Sentinel-1 images are created with Google Earth Engine (GEE) 111111https://earthengine.google.com directly in their corresponding projection and compressed with LZW (the standard compression in GEE). Sentinel-2 images are distributed via the Copernicus AWS via a .SAFE-archive format. We extract all needed bands in JPEG2000-format and build virtual raster files for combined access. JPEG2000 has the big advantage of being an extremely storage efficient, yet applies the compression on image-level, so windowed access is not possible.
@@
When doing inference, directly loading the entire tile and doing inference is not feasible, as data loading takes a long time due to the need to be decompressed. As this would lead to long GPU idle times, we use a different approach. We seperate the workload onto CPU- and GPU-nodes, where CPU-nodes load the data from the slow storage, decompress it multi-threaded and save it in a binary format on fast-access storage. The GPU-node then loads the fast binary and does inference. For each GPU-worker we start multiple CPU-workers to remove GPU idle time.
Appendix B Additional Figures

### v2

@@ Acknowledgements
This work was supported via the AI4Forest project, which is funded by the German Federal Ministry of Education and Research (BMBF; grant number 01IS23025A) and the French National Research Agency (ANR). It was also partially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy—The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689). We also acknowledge the computational resources provided by the PALMA II cluster at the University of Münster (subsidized by the DFG; INST 211/667-1) as well as by the Zuse Institute Berlin. We also appreciate the hardware donation of an A100 Tensor Core GPU from Nvidia and thank Google for their compute resources provided (Google Earth Engine).
@@
This study significantly advances forest monitoring by leveraging machine learning to incorporate temporal dynamics, resulting in the creation of tree canopy height maps over a four-year span (2019–2022) on a continental scale. Forests play a vital role in mitigating climate change by absorbing nearly half of human-generated carbon dioxide emissions (Friedlingstein et al., 2022). However, they are increasingly at risk due to climate change and human activities such as deforestation and land degradation (Anderegg et al., 2022). In support of global initiatives like the United Nations’ Sustainable Development Goals (SDGs)888https://sdgs.un.org/2030agenda, the Bonn Challenge999https://www.bonnchallenge.org, and the Glasgow Declaration101010https://www.oneplanetnetwork.org/programmes/sustainable-tourism/glasgow-declaration, effective forest management and conservation are crucial for climate adaptation and mitigation. Our approach significantly benefits the estimation of above-ground biomass, a critical component among the 55 essential climate variables (ECVs)111111https://gcos.wmo.int/site/global-climate-observing-system-gcos/essential-climate-variables/above-ground-biomass. A key improvement in this research is the enhanced detection of large trees, which is critical due to the power-law relationship between tree height and biomass. This advancement allows for more accurate biomass estimation, essential for assessing carbon stocks. Additionally, precise monitoring is crucial for validating carbon credit investments in the voluntary carbon market, ensuring the effectiveness and sustainability of forest growth projects while reducing potential leakage effects.
@@ Appendix A Data Handling Details
Handling big amounts of data is a challenge. To aid data handling and still having optimal projection accuracy, we downloaded, stored and preprocessed all data in the original Sentinel-2 tiling system. The tiling system is based on the UTM projections system but divides every UTM zone in tiles. Each UTM zone has its own projection121212For more details, we refer to: https://sentiwiki.copernicus.eu/web/s2-products.
Sentinel-1 images are created with Google Earth Engine (GEE) 131313https://earthengine.google.com directly in their corresponding projection and compressed with LZW (the standard compression in GEE). Sentinel-2 images are distributed via the Copernicus AWS via a .SAFE-archive format. We extract all needed bands in JPEG2000-format and build virtual raster files for combined access. JPEG2000 has the big advantage of being an extremely storage efficient, yet applies the compression on image-level, so windowed access is not possible.
@@
When doing inference, directly loading the entire tile and doing inference is not feasible, as data loading takes a long time due to the need to be decompressed. As this would lead to long GPU idle times, we use a different approach. We separate the workload onto CPU- and GPU-nodes, where CPU-nodes load the data from the slow storage, decompress it multi-threaded and save it in a binary format on fast-access storage. The GPU-node then loads the fast binary and does inference. For each GPU-worker we start multiple CPU-workers to remove GPU idle time.
Appendix B Ablation Studies
In the main paper we took several decisions without justification. Here we provide experimental ablation studies to underpin these conclusions.
B.1 Additional Sentinel-2 Bands
Sentinel-2 captures bands in resolutions of m, m and m and in this study we included all (except for B10, which is only for clouds). However, both m bands have wavelength, where direct knowledge transfer to this application is unknown: B01 measures coastal aerosols and B09 captures the density of water vapors. Our experiments show that both bands do not significantly increase the performance, neither do they decrease the performance. Further analysis is needed, but both bands are candidates to be removed from the set of input channels. Table 6 reports the results on the validation part of our dataset (L1 m refering to the L1 loss for all labels that exceed m)
| Configuration | L1 (m) | L1 m (m) | L1 m (m) | L1 m (m) | L1 m (m) | L2 (m) |
| Without B01 and B09 | 1.991 ± 0.002 | 4.837 ± 0.008 | 5.476 ± 0.010 | 7.384 ± 0.008 | 11.406 ± 0.004 | 22.281 ± 0.037 |
| Including B01 and B09 | 1.992 ± 0.003 | 4.830 ± 0.014 | 5.460 ± 0.015 | 7.364 ± 0.029 | 11.384 ± 0.031 | 22.277 ± 0.053 |
B.2 Time-Series: Seasonal pPatterns
Although our main paper shows that a time-series clearly outperforms the baseline setting of applying a median aggregation to the input data, we want to dive deeper in how the model leverages e.g. differences across the monthly images. We present 3 experiments: activation patterns for different months, performance for different forest types (with different phenological patterns) and an ablation between using only summer/winter months or both.
B.2.1 Activation patterns for different months
We use Guided Attention (Springenberg et al., 2014) to visualize the activation patterns across months for different patches (cf. Figure 7. We observe varying activation strengths across months and patches, suggesting the model processes temporal information differently by location. However, further research would be needed to confirm this hypothesis.
B.2.2 Performance on different forest types
| Model | Broadleaf | Coniferous |
| MAE (m) | MAE (m) | |
| Lang et al. | 5.44 | 5.11 |
| Liu et al. | 7.01 | 6.91 |
| Pauls et al. | 5.30 | 4.85 |
| Tolan et al. | 10.43 | 11.73 |
| Turubanova et al. | 8.60 | 8.43 |
| Ours | 4.57 | 4.11 |
We evaluated our model separately on broadleaf and coniferous forests using the Copernicus Land Monitoring Service Forest Type Map (2018) (Copernicus Land Monitoring Service, 2021). These forest types show different seasonal patterns - broadleaf forests have distinct leaf-on/off periods while coniferous forests maintain constant canopy. Table 7 reports our findings.
B.2.3 Seasonal information variation
To investigate the benefits of seasonal information independently from data volume (Note: the number of training labels remains identical for all variants), we conducted an ablation study comparing three models trained on different 4-month subsets: Winter (Nov-Feb), Summer (Jun-Sep) and Mixed (Jan-Feb, Aug-Sep).
| Model Variant | Huber Loss (m) |
| Winter (Nov–Feb) | 1.169 ± 0.003 |
| Summer (Jun–Sep) | 1.130 ± 0.002 |
| Mixed (Jan–Feb, Aug–Sep) | 1.122 ± 0.002 |
The results in Table 8 show that using only summer/leaf-on months is superior to using only winter/leaf-off months, however a mix of winter and summer months yields better validation performance than both individually.
Appendix C Analysis
Although our primary goal was to develop an openly available model that others can use for their own analyses, here we demonstrate the model’s temporal capabilities. We analyzed potential deforestation events by tracking pixels for which the corresponding height decreased from above 8m to below 5m between years. The affected area increased from between 2019 and 2020 and between 2020 and 2021 to between 2021 and 2022.
Appendix D Additional Figures